To get 90ml of a 35% sugar solution, a 65% solution is mixed with a 20% solution. How much of the 65% solution is needed?
Answer by stanbon(75887)

To get 90ml of a 35% sugar solution, a 65% solution is mixed with a 20% solution. How much of the 65% solution is needed?
-----
Equation:
sugar + sugar = sugar
0.65x + 0.20(90-x) = 0.35*90
---
Multiply thru by 100 to get:
65x + 20*90 - 20x = 35*90
45x = 15*90
x = (1/3)*90
x = 30ml (amt of 65% needed)
90-x = 60ml (amt of 20% needed)
===============================
Cheers,
Stan H.
================